# Gram-like matrix preserving extensions of noncommutative polynomials to sum of Hermitian Squares

### v1

@@ Acknowledgements
Both the authors extend their gratitude to Prof. Sumesh K and Prof. Arijit Dey for their suggestions regarding the article. The authors acknowledge Rajib Sarkar and Sreenanda S B for their assistance regarding the use of Macaulay2 software. The authors also thank Agnihotri Roy for reexamining some tedious calculations. The first named author would like to thank Indian Institute of Technology Madras for financial support (Office order No.F.ARU/R10/IPDF/2024). The second named author would like to thank Indian Institute of Technology Madras and ICSR for financial support (Project number - SB22231267MAETWO008573).

### v2

@@ Acknowledgements
Both the authors extend their gratitude to Prof. Sumesh K, Prof. Arijit Dey and Prof. Sriram B for their valuable suggestions regarding the article. The authors acknowledge Rajib Sarkar and Sreenanda S B for their assistance regarding the use of Macaulay2 software. The authors thank Dr. Aditi Howlader for suggesting some useful references. The authors also thank Agnihotri Roy for reexamining some tedious calculations. The first named author would like to thank Indian Institute of Technology Madras for financial support (Office order No.F.ARU/R10/IPDF/2024). The second named author would like to thank Indian Institute of Technology Madras and ICSR for financial support (Project number - SB22231267MAETWO008573).